OpenStudy (anonymous):

Solve. 2x-1/3+3<4 or 8x-2/2-1>6

OpenStudy (tkhunny):

Is it \(\dfrac{2x-1}{3}+3 < 4\) and \(\dfrac{8x-2}{2} - 1 > 6\)?

OpenStudy (anonymous):

yeah

OpenStudy (tkhunny):

That is very much NOT what you have written. Do you know why?

OpenStudy (anonymous):

Can you just show me how to solve it .

OpenStudy (tkhunny):

You don't seem to have answered my question. Learning to comunicate it very important.

OpenStudy (anonymous):

yes i know why, just didnt feel like writing correctly, so can you just show me how to solve it thats why im asking for help.

OpenStudy (tkhunny):

You're also going to have to learn to cooperate, better. Tell me how the notation could be improved.

OpenStudy (anonymous):

if your not going to help me, i prefer you leave because i didnt ask for a lesson im asking for an explanation.

OpenStudy (tkhunny):

Fair enough. The very first thing you need to do to solve this problem is to write the pieces clearly and communicate them accurately. Please do step #1 and we can continue.

OpenStudy (tkhunny):

(2x-1)/3 + 3 < 4 Subtract 3 (2x-1)/3 + 3 - 3 < 4 - 3 Simplify (2x-1)/3 < 1 Multiply by 3 [(2x-1)/3]*3 < 1*3 Simplify 2x-1 < 3 Add 1 2x-1+1 < 3+1 Simplify 2x < 4 Divide by 2 2x/2 < 4/2 Simplify x < 2 You do the other one.

OpenStudy (anonymous):

is it x>2 ?

OpenStudy (tkhunny):

I wrote "x<2" and you asked "x>2"? Why?

OpenStudy (anonymous):

Because thats the answer to the other one

OpenStudy (tkhunny):

Did you show ANY of your work? Seriously, you must learn to COMMUNCATE. This is your mathematics lesson for the day. It has EVERYTHING to do with inequalities and algebraic manipulation. If you cannot communicate your work, it is all of no value.

OpenStudy (tkhunny):

Okay, you have, after a little algebra, x < 2 or x > 2 What, then, is your final solution and in what form shall we write it so that it can be communicated clearly and completely?

OpenStudy (anonymous):

nevermind i dont need to solve that inequality anymore i need to solve this: \[-\frac{ 2 }{ 5}x-9<\frac{ 9 }{ 10 }\]

OpenStudy (anonymous):

can you solve it, with an explanation.

OpenStudy (tkhunny):

Yes, I can. Can you show any work at all?